# Exact Subgraph Isomorphism Network for Predictive Graph Mining

### v1

@@ 2 $H_{0}\leftarrow\text{a graph at root of mining tree}$
5       ${\cal W}\leftarrow\mathrm{Traverse}(H_{0},{\cal W},\lambda)$
6         Update $\boldsymbol{B}$ by (4) for $H\in{\cal W}$
7         Update $\boldsymbol{b}$ by (5)
8       for iter = 1, …, MaxIter do
9            Update $\boldsymbol{\Theta}$ by (6)
10
11         end for
12
@@ 9            Update $\boldsymbol{\Theta}$ by (6)
1function Traverse($H,{\cal W},\lambda$)
2 if $\mathrm{UB}(H)\leq\lambda$ then return $\cal W$}
3 if $\|\boldsymbol{g}_{H}\|_{2}>\lambda$ then ${\cal W}\leftarrow{\cal W}\cup\{H\}$
4 ${\cal C}\leftarrow\mathrm{Expand}(H)$
5 for $H^{\prime}\in{\cal C}$ do
6    ${\cal W}\leftarrow\mathrm{Traverse}(H^{\prime},{\cal W},\lambda)$
@@
13    $\cal C$$\leftarrow$ all graphs expanded from $H$ by gSpan
14      Set $\{\psi_{H^{\prime}}(G_{i})\}_{i=1}^{n}$ for $H^{\prime}\in{\cal C}$
15
@@ 14      Set $\{\psi_{H^{\prime}}(G_{i})\}_{i=1}^{n}$ for $H^{\prime}\in{\cal C}$
17    $\cal C$$\leftarrow$ Retrieve already expanded children of $H$
18
@@ 17    $\cal C$$\leftarrow$ Retrieve already expanded children of $H$
22
Algorithm 2 Working Set Generation
By combining the rule (10) and a graph mining algorithm, we can build an efficient pruning algorithm. We employ a well-known graph mining algorithm called gSpan (Yan and Han,, 2002) that can efficiently enumerate all the subgraphs in a given set of graphs. Figure 1 (c) is an illustration of a graph mining tree. At each tree node, gSpan expands the graph (add an edge and a node) as far as the expanded graph is included in the given dataset as a subgraph. By providing a unique code to each generated graphs (called the DFS code because it is based on depth-first search in a graph), gSpan can enumerate all the subgraphs without generating duplicated graphs (For further detail, see (Yan and Han,, 2002)). As shown in Fig. 1 (c), all the graphs in the tree contains their ancestors as a subgraph. As a result, the following important consequence is obtained
@@
[…]
@@
Since our pruning strategy only omit the update of unnecessarily parameters, i.e., $\boldsymbol{\beta}_{H}=\boldsymbol{0}$, the pruning does no change all the parameter values compared with the case that we do not use the pruning. This also means that the final prediction performance also does not change because of the pruning.
@@ 2.2.2 Algorithm
We here describe the optimization procedure of EIN. When the regularization parameter $\lambda$ is large, $\boldsymbol{B}$ becomes highly sparse, by which computations usually become faster because more subgraphs are expected to be pruned. Therefore, after starting from a larger value of $\lambda$, we gradually reduce $\lambda$ while optimizing parameters. Let $\Lambda=(\lambda_{1},\ldots,\lambda_{K})$ be a sequence of the regularization parameters, where $\lambda_{1}>\lambda_{2}\cdots>\lambda_{K}$. For each $\lambda$, we use the solution of previous $\lambda$ as an initial solution.
@@
The Train-EIN function of Algorithm 1 performs the optimization (2) for each $\lambda$ from the given $\Lambda$. In line 5, the Traverse function performs the graph mining tree search in which $\cal W$} is generated via gradient pruning. The detailed procedure of the working set generation is in Algorithm 2. The Traverse function searches graph mining tree recursively. At each tree node, $\mathrm{UB}(H)$ is evaluated and the entire subtree below the current node can be pruned if $\mathrm{UB}(H)\leq\lambda$. On the other hand, if $\|\boldsymbol{g}_{H}\|_{2}>\lambda$, $H$ should be included in $\cal W$}. The children nodes of $H$ are created by gSpan in the Expand function, by which only the subgraphs $H^{\prime}$ included in the training dataset can be generated and as a byproduct of this process, we obtain $\psi_{H^{\prime}}(G_{i})$. Note that if the children are already generated in the previous iterations, we can reuse them.
Once $\cal W$} is determined, $\boldsymbol{B}$ and $\boldsymbol{b}$ are updated in line 6-7 in Algorithm 1. Their step lengths are determined by backtrack algorithm (Note that since our pruning condition does not depend on the step length, we do not need to perform the Traverse function repeatedly during this backtrack steps). Then, $\boldsymbol{\Theta}$ is updated in line 8-10. This is a usual neural network parameter update, for which we iterate the update until given max iteration. Here, we employed the full instance gradient, the standard stochastic gradient can also be used. At each $\lambda$, the alternating update procedure stops (line 11) when the iteration reaches the given maximum iterations or some other stopping condition is satisfied.
@@ 2.3 Post-hoc Analysis for Knowledge Discovery
After the optimization, only a small number of $\boldsymbol{\beta}_{H}$ have non-zero values, by which we can identify predictive subgraphs. Let ${\cal S}=\{H\in{\cal H}\mid\boldsymbol{\beta}_{H}\neq\boldsymbol{0}\}$ be the set of the selected subgraphs by EIN. In our later experiments, we observed that $|{\cal S}|$ was typically ranged from around 50 to at most a few hundreds. Then, some insight may be obtained just by directly observing all of those selected subgraphs by the domain expert of the data.
The trained EIN can be seen as an $|{\cal S}|$-dimensional input neural network, to which post-hoc knowledge discovery methods (Bhati et al.,, 2025) can be applied. For example, well-known SHAP (Lundberg and Lee,, 2017) and LIME (Ribeiro et al.,, 2016), both of which provide local feature importance, is applicable. Another typical approach is to use the trained EIN as a teacher and fit an interpretable surrogate model (e.g., decision tree), from which a possible decision rule can be estimated (Molnar,, 2025). Note that, usually, applying interpretable machine learning methods directly to exhaustive subgraph isomorphism features can be computationally intractable. Because of EIN, which selects a few important subgraphs, a variety of post-hoc analyses become much easier.
@@ 2.4 Combining with Graph Neural Network
where $\oplus$ is the vector concatenation, $\mathrm{GNN}$ is any GNN that outputs arbitrary dimension representation vector, and $\boldsymbol{\Theta}_{\rm GNN}$ and $\boldsymbol{\Theta}_{\rm FFN}$ are parameters of $\mathrm{GNN}$ and $\mathrm{FFN}$, respectively. This combination enables any GNNs to enhance the discriminative ability in terms of the selected subgraphs by EIN. A schematic illustration of a simplest case is shown in Fig. 2. Computations of the combined model can also be performed by almost the same alternating update (4)-(6). The only difference is that the FFN parameter update (6) is replaced with the simultaneous update of $\boldsymbol{\Theta}_{\rm FFN}$ and $\boldsymbol{\Theta}_{\rm GNN}$.
@@ 3 Related Work
For GNNs, the message passing based approach has been widely employed (Zhou et al.,, 2020). It is known that the expressive power of classical message passing approaches is limited by the first order Weisfeiler-Lehman (1-WL) test (Zhao et al.,, 2022). Many studies have tackled this limitation, some of which actually can be comparable with higher order WL tests. For example, $k$-hop extensions of the message passing (e.g., Abu-El-Haija et al.,, 2019; Nikolentzos et al.,, 2020; Wang et al.,, 2021; Chien et al.,, 2021; Brossard et al.,, 2020) and the higher order (tensor) representations (e.g., Morris et al.,, 2019; Maron et al., 2019b, ; Maron et al., 2019a, ; Maron et al., 2019c, ; Keriven and Peyré,, 2019; Geerts and Reutter,, 2022; Murphy et al.,, 2019) are popular approaches. Further, several studies (Bouritsas et al.,, 2023; Cotta et al.,, 2021; Barcelo et al.,, 2021; Bevilacqua et al.,, 2022) incorporate small subgraph information, such as pre-specified motifs and ego-networks, into neural networks. On the other hand, EIN performs the data-adaptive selection of a small number of (globally) predictive subgraphs, based on the exact enumeration of all the subgraphs in the dataset, which is obviously different approach from the above popular GNN studies. Further, most of GNNs can be combined with EIN by (11).
@@
In the context of graph mining, discriminative pattern mining has also been studied (e.g., Thoma et al.,, 2010; Potin et al.,, 2025; Chen et al.,, 2022). However, most of them are not directly incorporated in a learning model. Typically, the subgraph enumeration (using some pruning strategies) is performed based on a discriminative score (e.g., ratio of frequency between two classes) independent from a learning model, and the selection results can be used in a subsequent prediction model training. Although (Nakagawa et al.,, 2016; Yoshida et al.,, 2019, 2021; Tajima et al.,, 2024) enumerate subgraphs during a model optimization, they are limited to a linear combination of the subgraphs. Note that Tajima et al., (2024) also use a combination of proximal gradient and gSpan, from which our optimization procedure is extended to the neural network backpropagation and the group-sparsity. Although (Nakagawa et al.,, 2016; Yoshida et al.,, 2019, 2021) provide a stronger pruning rule that can satisfy the global optimality, the convex formulation of the model training is required. To our knowledge, a discriminative subgraph mining that directly combines the exact subgraph enumeration with a (sparsity-induced) neural network has not been investigated.

### v2

@@ 4 $H_{0}\leftarrow\text{a graph at root of mining tree}$
7         Calculate $\eta_{H}$ for $\forall H\in S_{t}$
8       $S_{t+\frac{1}{2}}\leftarrow S_{t}$
9       $\zeta\leftarrow s\operatorname{-max}\left\{\eta_{H}\right\}_{H\in S_{t+\frac{1}{2}}}$
10       $S_{t+\frac{1}{2}}\leftarrow\mathrm{Traverse}(H_{0},s,\zeta,S_{t+\frac{1}{2}})$
11       $\boldsymbol{B}^{t+1}_{S_{t+\frac{1}{2}}}\leftarrow H_{s}(\boldsymbol{B}_{S_{t+\frac{1}{2}}}^{t}-\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{B}_{S_{t+\frac{1}{2}}}}\ell^{t})$, where $\gamma_{\boldsymbol{B}}^{t}$ is determined by backtrack search
12       $\boldsymbol{b}^{t+1}\leftarrow\mathrm{GradientDescent}(\boldsymbol{b}^{t})$
13       $\boldsymbol{\theta}^{t+1}\leftarrow\mathrm{GradientDescent}(\boldsymbol{\theta}^{t})$
@@
1function Traverse($H,s,\zeta,S_{t+\frac{1}{2}}$)
2   Calculate $\eta_{H}$ and $\mathrm{UB}(H)$ for current $H$
3 if $\mathrm{UB}(H)<\zeta$ then
    return $S_{t+\frac{1}{2}}$
[…]
4
5 else if $\eta_{H}>\zeta$ then
6    $S_{t+\frac{1}{2}}\leftarrow S_{t+\frac{1}{2}}\cup\{H\}$
7    $\zeta=s\operatorname{-max}\left\{\eta_{H^{\prime}}\right\}_{H^{\prime}\in S_{t+\frac{1}{2}}}$
8    $S_{t+\frac{1}{2}}\leftarrow S_{t+\frac{1}{2}}\setminus\{H^{\prime}\mid\eta_{H^{\prime}}<\zeta,\ H^{\prime}\notin S_{t}\}$
9   end if
10 ${\mathcal{C}}\leftarrow\mathrm{Expand}(H)$
11 for $H^{\prime}\in{\mathcal{C}}$ do
12    $S_{t+\frac{1}{2}}\leftarrow\mathrm{Traverse}(H,s,\zeta,S_{t+\frac{1}{2}})$
@@
19    $\mathcal{C}$$\leftarrow$ all graphs expanded from $H$ by gSpan
20      Set $\{\psi_{H^{\prime}}(G_{i})\}_{i=1}^{n}$ for $H^{\prime}\in{\mathcal{C}}$
21
@@ 20      Set $\{\psi_{H^{\prime}}(G_{i})\}_{i=1}^{n}$ for $H^{\prime}\in{\mathcal{C}}$
23    $\mathcal{C}$$\leftarrow$ Retrieve already expanded children of $H$
24
@@ 23    $\mathcal{C}$$\leftarrow$ Retrieve already expanded children of $H$
28
Algorithm 2 Traverse Mining Tree
As a result, the following important consequence is obtained
@@
By evaluating $\mathrm{UB}(H)\leq\zeta$ during the tree traverse of gSpan (e.g., depth-first search), all the descendants of $H$ in the tree can be pruned if the inequality holds. This means that we can perform the update (3) without enumerating all the possible subgraphs $\mathcal{H}$}.
@@
Since our pruning strategy only omit the update of unnecessarily parameters, i.e., $\boldsymbol{\beta}_{H}=\boldsymbol{0}$, the pruning does no change any parameter values compared with the case that we do not use the pruning. This also means that the final prediction performance also does not change because of the pruning.
@@ 2.2.2 Algorithm
We here describe the optimization procedure of EIN. The basic strategy is that we first set $s$ as a small value (e.g., $1$) and gradually increase $s$ while optimizing parameters. Let ${\mathcal{S}}=(s_{1},\ldots,s_{K})$ be a sequence of $s$, where $s_{1}<s_{2}<\cdots<s_{K}$. For each $s$, we use the solution of previous $s$ as an initial solution.
@@
The Train-EIN function of Algorithm 1 performs the optimization (2) for each $s$. In line 10, the Traverse function performs the graph mining tree search that generates a set of candidate subgraphs denoted as $S_{t+\frac{1}{2}}$, consisting of subgraphs that are not pruned (i.e., $S_{t+1}\subseteq S_{t+\frac{1}{2}}$). The detailed procedure of the candidate generation is in Algorithm 2. The Traverse function recursively searches the graph mining tree. At each tree node, $\mathrm{UB}(H)$ is evaluated and the entire subtree below the current node can be pruned if $\mathrm{UB}(H)<\zeta$ (line 3-4). On the other hand, if $\eta_{H}>\zeta$, $H$ is added to $S_{t+\frac{1}{2}}$ (line 6). Then, the threshold $\zeta$ is updated (line 7), after which we can remove $\{H^{\prime}\mid\eta_{H^{\prime}}<\zeta,\ H^{\prime}\notin S_{t}\}$ (line 8). This guarantees $|S_{t+\frac{1}{2}}|\leq 2s$ (for simplicity, we assume there is no tie of $\eta_{H}$). The children nodes of $H$ are created by gSpan in the Expand function, by which only the subgraphs $H^{\prime}$ included in the training dataset can be generated and as a byproduct of this process, we obtain $\psi_{H^{\prime}}(G_{i})$. Note that if the children are already generated in the previous iterations, we can reuse them. Once $S_{t+\frac{1}{2}}$ is determined, $\boldsymbol{B}^{t}$ is updated in line 11 of Algorithm 1, in which the step length $\gamma_{\boldsymbol{B}}^{t}$ is determined by the batcktrack search. Here, $\boldsymbol{B}_{S_{t+\frac{1}{2}}}$ is a submatrix of $\boldsymbol{B}$ in which only the columns specified by $S_{t+\frac{1}{2}}$ are extracted. In line 12-13 of Algorithm 1, $\boldsymbol{b}$ and $\boldsymbol{\theta}$ are updated by the standard gradient descent, in which the batcktrack search is also performed.
2.2.3 Convergence Analysis
We here show that EIN can converge to a critical point with at least sublinear convergence.
Suppose that the gradient descent of $\boldsymbol{b}$ and $\boldsymbol{\theta}$ iterates $\tau_{\max}$ times. The index $\tau$ represents the number of iteration of this inner loop as follows.
$\boldsymbol{b}^{t,\tau+1}$ $=\boldsymbol{b}^{t,\tau}-\gamma^{t,\tau}_{\boldsymbol{b}}\nabla_{\boldsymbol{b}}\ell^{t,\tau},$
$\boldsymbol{\theta}^{t,\tau+1}$ $=\boldsymbol{\theta}^{t,\tau}-\gamma^{t,\tau}_{\boldsymbol{\theta}}\nabla_{\boldsymbol{\theta}}\ell^{t,\tau},$
where $\nabla_{\boldsymbol{b}}\ell^{t,\tau}={\frac{\partial\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t,\tau},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{b}}}$, $\nabla_{\boldsymbol{\theta}}\ell^{t,\tau}={\frac{\partial\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t,\tau})}{\partial\boldsymbol{\theta}}}$, and $\gamma^{t,\tau}_{\boldsymbol{b}}$ and $\gamma^{t,\tau}_{\boldsymbol{\theta}}$ are the step lengths of the $\tau$-th iteration, respectively. Let $c_{\boldsymbol{B}},c_{\boldsymbol{b}},c_{\boldsymbol{\theta}}\in(0,1)$ be constants used in the backtrack termination conditions. Our backtrack, in which the step length iteratively decays with a ratio $\rho\in(0,1)$, stops when
$\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})$ $\geq{c_{\boldsymbol{B}}(1-\rho)\frac{\gamma^{t}_{\boldsymbol{B}}}{2}\|\nabla_{\boldsymbol{B}}\ell^{t}_{S}\|_{F}^{2}}$ (7)
$\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t,\tau},\boldsymbol{\theta}^{t})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t,\tau+1},\boldsymbol{\theta}^{t})$ $\geq c_{\boldsymbol{b}}\frac{\gamma^{t,\tau}_{\boldsymbol{b}}}{2}\|\nabla_{\boldsymbol{b}}\ell^{t,\tau}\|_{2}^{2}$ (8)
$\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t,\tau})-\ell(\boldsymbol{B}^{t+1},\boldsymbol{b}^{t+1},\boldsymbol{\theta}^{t,\tau+1})$ $\geq c_{\boldsymbol{\theta}}\frac{\gamma^{t,\tau}_{\boldsymbol{\theta}}}{2}\|\nabla_{\boldsymbol{\theta}}\ell^{t,\tau}\|_{2}^{2}$ (9)
are satisfied, respectively. These conditions are from the well-known Armijo condition (Nocedal and Wright,, 2006).
Throughout the proof, we assume differentiability and $L$-smoothness of $\ell$:
Definition 2.1.
Let $L$ be the Lipschitz constant and $\boldsymbol{z}$ be the vector that concatenates all the elements of $\boldsymbol{B}$, $\boldsymbol{b}$, and $\boldsymbol{\theta}$. The (restricted) $L$-smoothness of $\ell$ is defined as
$\left\|{\frac{\partial\ell(\boldsymbol{B},\boldsymbol{b},\boldsymbol{\theta})}{\partial\boldsymbol{z}}}-{\frac{\partial\ell(\boldsymbol{B}^{\prime},\boldsymbol{b}^{\prime},\boldsymbol{\theta}^{\prime})}{\partial\boldsymbol{z}}}\right\|_{2}\leq L\|\boldsymbol{z}-\boldsymbol{z}^{\prime}\|_{2}$ (10)
for any pair of $(\boldsymbol{B},\boldsymbol{b},\boldsymbol{\theta})$ and $(\boldsymbol{B}^{\prime},\boldsymbol{b}^{\prime},\boldsymbol{\theta}^{\prime})$, where $\|\boldsymbol{B}\|_{0}\leq s$, $\|\boldsymbol{B}^{\prime}\|_{0}\leq s$.
We first guarantee the existence of the positive step lengths that satisfy the backtrack conditions.
Lemma 2.1.
There exist positive step lengths $\gamma_{\boldsymbol{B}}^{t}\geq\rho^{2}/L$, $\gamma_{\boldsymbol{b}}^{t,\tau}\geq\rho/L$, and $\gamma_{\boldsymbol{\theta}}^{t,\tau}\geq\rho/L$ that satisfy (7), (8), and (9), respectively.
The proof is in Appendix E. For $\boldsymbol{B}$, the proof is mainly from (Damadi and Shen,, 2022), and for $\boldsymbol{b}$ and $\boldsymbol{\theta}$, it is from a known property of the gradient descent. The convergence of our procedure is shown by the following theorem, in which our problem (2) is re-written as an equivalent unconstrained problem:
Theorem 2.2.
Let
$\tilde{\ell}(\boldsymbol{z})=\ell(\boldsymbol{B},\boldsymbol{b},\boldsymbol{\theta})+\delta_{s}(\boldsymbol{B})$
be the objective function with the penalty on the constraint violation
$\delta_{s}(\boldsymbol{B})=\begin{cases}0,&\text{ if }\|\boldsymbol{B}\|_{2,0}\leq s,\\ \infty,&\text{ otherwise. }\end{cases}$
The sequence $\boldsymbol{z}^{t}$, defined by $\boldsymbol{B}^{t}$, $\boldsymbol{b}^{t}$, and $\boldsymbol{\theta}^{t}$, is assumed to be bounded. Then, $\boldsymbol{z}^{t}$ is a Cauchy sequence that converges to a critical point $\boldsymbol{z}^{*}$ satisfying $\partial\tilde{\ell}(\boldsymbol{z}^{*})\ni\boldsymbol{0}$, where $\partial\tilde{\ell}(\boldsymbol{z})$ is the limiting subdifferential of $\tilde{\ell}$.
See Appendix G for the proof. The limiting subdifferential (Mordukhovich,, 2012) is a generalized notion of subdifferential applicable to an arbitrary proper and lower semi-continuous (extended-real-valued) function (See Appendix F for the definition). The generalized Fermat’s rule (Rockafellar et al.,, 2009) indicates $\partial\tilde{\ell}(\boldsymbol{z}^{*})\ni\boldsymbol{0}$ is a necessarily condition of a local minimum. The boundedness assumption of $\boldsymbol{z}^{t}$ is introduced to ensure that $\boldsymbol{z}^{t}$ does not diverge to infinity. In practice, if one wishes to guarantee this rigorously, a possible approach is to introduce an $L_{2}$ penalty (weight decay) on the parameters (for which the regularization coefficient can be chosen to be arbitrarily small). We further derive the following convergence rate:
Theorem 2.3.
Under the same assumption as Theorem 2.2, our optimization algorithm has at least the following sublinear convergence rate:
$\|\boldsymbol{z}^{t}-\boldsymbol{z}^{*}\|_{2}=O(t^{q}),$
where $q<0$.
Based on properties of $\boldsymbol{z}^{t}$ obtained through the proof of Theorem 2.2, the above rate can be derived by the almost same proof as Theorem 2 of (Attouch and Bolte,, 2009). We show the proof in Appendix H for completeness.
@@ 2.3 Post-hoc Analysis for Knowledge Discovery
After the optimization, only $s$ subgraphs have non-zero values of $\boldsymbol{\beta}_{H}$, by which we can identify predictive subgraphs. In our later experiments, we observed that small $s$ (up to $100$) can achieve sufficiently high predictive performance. Therefore, examining the selected subgraphs—possibly by a domain expert—may lead to new insights about the data.
The trained EIN can be seen as an $s$-dimensional input neural network, to which post-hoc knowledge discovery methods (Bhati et al.,, 2025) can be applied. For example, well-known SHAP (Lundberg and Lee,, 2017) and LIME (Ribeiro et al.,, 2016), both of which provide local feature importance, is applicable. Another typical approach is to use the trained EIN as a teacher and fit an interpretable surrogate model (e.g., decision tree), from which a possible decision rule can be estimated (Molnar,, 2025). Note that, usually, applying interpretable machine learning methods directly to exhaustive subgraph features can be computationally intractable. Benefiting from EIN, we can only focus on a small number of important selected subgraphs, by which a variety of post-hoc analyses become much easier to apply.
@@ 2.4 Combining with Graph Neural Network
where $\oplus$ is the vector concatenation, $\mathrm{GNN}$ is any GNN that outputs arbitrary dimension representation vector, and $\boldsymbol{\theta}_{\rm GNN}$ and $\boldsymbol{\theta}_{\rm FFN}$ are parameters of $\mathrm{GNN}$ and $\mathrm{FFN}$, respectively. This combination enables any GNNs to enhance the discriminative ability in terms of the selected subgraphs by EIN. A schematic illustration of a simplest case is shown in Fig. 2. Computations of the combined model can also be performed by almost the same alternating update shown in Algorithm 1. The only difference is that the FFN parameter update (line 13 of Algorithm 1) is replaced with the simultaneous update of $\boldsymbol{\theta}_{\rm FFN}$ and $\boldsymbol{\theta}_{\rm GNN}$.
@@ 3 Related Work
For GNNs, the message passing based approach has been widely employed (Zhou et al.,, 2020). It is known that the expressive power of classical message passing approaches is limited by the first order Weisfeiler-Lehman (1-WL) test (Zhao et al.,, 2022). Many studies have tackled this limitation, some of which actually can be comparable with higher order WL tests. For example, $k$-hop extensions of the message passing (e.g., Abu-El-Haija et al.,, 2019; Nikolentzos et al.,, 2020; Wang et al.,, 2021; Chien et al.,, 2021; Brossard et al.,, 2020) and the higher order (tensor) representations (e.g., Morris et al.,, 2019; Maron et al., 2019b, ; Maron et al., 2019a, ; Maron et al., 2019c, ; Keriven and Peyré,, 2019; Geerts and Reutter,, 2022; Murphy et al.,, 2019) are popular approaches. Further, several studies (Bouritsas et al.,, 2023; Cotta et al.,, 2021; Barcelo et al.,, 2021; Bevilacqua et al.,, 2022; Yan et al.,, 2024) incorporate small subgraph information, such as pre-specified motifs and ego-networks, into neural networks. On the other hand, EIN performs the data-adaptive selection of a small number of (globally) predictive subgraphs, based on the exact enumeration of all the subgraphs in the dataset, which is obviously different approach from the above popular GNN studies. Further, most of GNNs can be combined with EIN by (11).
@@
In the context of graph mining, discriminative pattern mining has also been studied (e.g., Thoma et al.,, 2010; Potin et al.,, 2025; Chen et al.,, 2022). However, most of them are not directly incorporated in a learning model. Typically, the subgraph enumeration (using some pruning strategies) is performed based on a discriminative score (e.g., ratio of frequency between two classes) independent from a learning model, and the selection results can be used in a subsequent prediction model training. Although (Nakagawa et al.,, 2016; Yoshida et al.,, 2019, 2021; Tajima et al.,, 2024) enumerate subgraphs during a model optimization, they are limited to a linear combination of the subgraphs. Although (Nakagawa et al.,, 2016; Yoshida et al.,, 2019, 2021) provide a stronger pruning rule that can satisfy the global optimality, the convex formulation of the model training is required. Several studies have considered learning $L_{0}$ based sparse neural networks (e.g., Louizos et al.,, 2018; Damadi et al.,, 2024). However, to our knowledge, a discriminative subgraph mining that directly combines the exact subgraph enumeration with an $L_{0,2}$ constraint neural network has not been investigated.